# NLP-enabled trajectory map-matching in urban road networks using transformer sequence-to-sequence model

### v1

@@ Abstract
Large-scale geolocation telematics data acquired from connected vehicles has the potential to significantly enhance mobility infrastructures and operational systems within smart cities. To effectively utilize this data, it is essential to accurately match the geolocation data to the road segments. However, this matching is often not trivial due to the low sampling rate and errors exacerbated by multipath effects in urban environments. Traditionally, statistical modeling techniques such as Hidden-Markov models incorporating domain knowledge into the matching process have been extensively used for map-matching tasks. However, rule-based map-matching tasks are noise-sensitive and inefficient in processing large-scale trajectory data. Deep learning techniques directly learn the relationship between observed data and road networks from the data, often without the need for hand-crafted rules or domain knowledge. This renders them an efficient approach for map-matching large-scale datasets and more robust to the noise. This paper introduces a sequence-to-sequence deep-learning model, specifically the transformer-based encoder-decoder model, to perform as a surrogate for map-matching algorithms. The encoder-decoder architecture initially encodes the series of noisy GPS points into a representation that automatically captures autoregressive behavior and spatial correlations between GPS points. Subsequently, the decoder associates data points with the road network features and thus transforms these representations into a sequence of road segments. The model is trained and evaluated using GPS traces collected in Manhattan, New York. Achieving an accuracy of 76%, transformer-based encoder-decoder models extensively employed in natural language processing presented a promising performance for translating noisy GPS data to the navigated routes in urban road networks.
@@ III-A Map-matching preliminaries
Map-matching is a technique designed to determine the sequence of connected road segments on a digital road network map that best approximates the actual route taken by a vehicle, even in the presence of errors and noise in the observed GPS data. This task is challenging, especially in an urban environment, due to the significant positioning errors and the complex, dense road map. In our study, the map-matching algorithm takes a sequence of observed coordinates in raw GPS data as input and produces a segment-based trajectory as the output. However, the low sampling rate in data collection leads to a sparse input representation. Consequently, no direct one-to-one correspondence exists between observed GPS points and inferred road segments.
@@ IV-C Objective Function
Let $g=(g_{1},g_{2},...,g_{n})$ be the transformed version of $T=\left\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})\right\}$, representing the variable-length input sequence, here in this study, a sequence of grid cells, each of which encompasses each GPS point, and $r=(r_{1},r_{2},...,r_{m})$ be the variable-length output sequence, i.e., sequence of road segments, $m$ is the length of the output sequence. The seq2seq model maps $g$ to $r$ by learning a conditional probability distribution over the set of possible output sequences, given the input sequence.
$p(r|g)=\prod_{t=1}^{m}p(r_{t}|r_{1},\dots,r_{t-1},g)$
The model learns a probability distribution from input-output pairs in training, adjusting parameters to maximize output sequence likelihood given input; the seq2seq model aims to maximize this log-likelihood.
$arg\max\limits_{\theta}\ P(r|g)=argmax\log P(r|g)\\ =argmax\sum_{t=1}^{m}logP(r_{t}|g_{n},r_{<t}).$
@@ IV-D1 Encoder
$PE_{(pos,i)}=sin(\frac{pos}{10000^{(\frac{2i}{d_{emb}})}}),\$ $PE_{(pos,2i+1)}=cos(\frac{pos}{10000^{(\frac{2i}{d_{emb}})}}),\$ where $pos\in\{1,\dots,n\}$ is the position of the token in the sequence of length $n$, $i\in\{1,\dots,d_{emb}\}$ is the dimension of the positional encoding, $d_{emb}$ is the model’s embedding dimension also referred as model dimension.
@@ IV-D2 Decoder
V Experiment
V-A Baseline
V-A1 HMM
@@
V-A2 RNN-based encoder-decoder with attention mechanism
@@
The main components of the RNN-based encoder-decoder architecture are two layers: the embedding layer and the RNN layer. Therefore, the main difference between the RNN-based encoder-decoder with the transformer-based encoder-decoder is its recurrent layer versus the self-attention layer in the transformer. The architecture of the RNN-based encoder-decoder model is presented in Figure 3.
Figure 3: Bi-GRU Encoder-Decoder
In an RNN-based encoder-decoder, during the encoding, the RNN layer processes all the tokens on the input sequence iteratively and reutilizes the hidden state of an iteration as the input for generating the hidden state for the next element. The last hidden state of the RNN is passed as input to the first iteration of the decoder. Therefore, when processing a sequence of inputs, the RNN layer generates a sequence of hidden states, with each hidden state being a function of the current input and the previous hidden state. In particular, a bidirectional Gated Recurrent Unit (Bi-GRU) in the RNN layer of the encoder, and a regular GRU in the RNN layer of the decoder are utilized. The structure of the Gated Recurrent Unit (GRU) is presented in figure 4.
Figure 4: Gated Recurrent Unit
GRU captures the sequential information in the data by maintaining an internal hidden state that evolves over time. It has a gating mechanism that controls the flow of information within the network, making it capable of learning and remembering longer sequences compared to traditional RNNs. The gating mechanism consists of an update gate and a reset gate, which control how much of the previous hidden state should be retained and how much new information should be incorporated in each time step.
A Bi-GRU is a variant of the GRU architecture that processes input sequences in both forward and backward directions. The input sequence is fed into two separate GRU layers, one processing the input sequence in the forward direction and the other processing the input sequence in the backward direction. The outputs of the two layers at each time step are concatenated and passed to the next layer or used as the final output. The architecture of the GRU cell is shown in figure 4.
The output of each direction of GRU, $h_{t}$, called hidden state at time step $t$, is a linear combination of the previous hidden state $h_{t-1}$ and candidate activation $\tilde{h}_{t}$ in which an update vector, $u_{t}$, controls the extent of the combination. The candidate activation $\tilde{h}_{t}$ is a function of the input, $x_{t}$, and previous hidden state after going through the reset gates, and similarly, the update vector, $u_{t}$, is a function of the input and previous hidden state after going through the update gates. The equations that describe the GRU update process for forward and backward GRU are as follows. The superscripts $f$ and $b$ denote the forward and backward recurrent components, respectively.
$h_{t}^{f}=(1-u_{t}^{f})*h_{t-1}^{f}+u_{t}^{f}*\tilde{h}_{t}^{f},\\ \tilde{h}_{t}^{f}=\phi(W_{h}^{f}[r_{t}^{f}*h_{t-1}^{f},x_{t}]+b_{h}),\\ u_{t}^{f}=\sigma(W_{u}^{f}[h_{t-1}^{f},x_{t}]+b_{u}^{f}),\\ r_{t}^{f}=\sigma(W_{r}^{f}[h_{t-1}^{f},x_{t}]+b_{r}^{f}),\$
$h_{t}^{b}=(1-u_{t}^{b})*h_{t+1}^{b}+u_{t}^{b}*\tilde{h}_{t}^{b}\\ \tilde{h}_{t}^{b}=\phi(W_{h}^{b}[r_{t}^{b}*h_{t+1}^{b},x_{t}]+b_{h}^{b}),\\ r_{t}^{b}=\sigma(W_{r}^{b}[h_{t+1}^{b},x_{t}]+b_{r}^{b}),\\ u_{t}^{b}=\sigma(W_{u}^{b}[h_{t+1}^{b},x_{t}]+b_{u}^{b}),\$
where $r_{t}^{f}$ and $r_{t}^{b}$ are reset vectors, $u_{t}^{f}$ and $u_{t}^{b}$ are update vectors, $\tilde{h}_{t}^{f}$ and $\tilde{h}_{t}^{b}$ are candidate activations, $h_{t}^{f}$ and $h_{t}^{b}$ are the hidden states. $x_{t}$ is the current input, and $t$ denotes the time step in the input sequence. $\sigma$ and $\phi$ are the sigmoid and hyperbolic tangent activation functions, respectively. $W_{h}$, $W_{r}$, $W_{z}$, $w_{y}$ are weight matrices associated with the input vectors and $b_{h}$, $b_{u}$, $b_{r}$ are the biases which are the model’s learnable parameters. The hidden states of the two forward and backward layers at each time step are concatenated to produce the final activation state, $h_{t}=[h_{t}^{f},h_{t}^{b}]$.
In a regular encoder-decoder setup, a single hidden state is used to represent the entire input sequence and convey all the information from the encoder to the decoder. However, relying solely on this single vector can result in a loss of important information over time, as the information gets diluted by the time it reaches the final state of the encoder. To address this issue, an attention mechanism is used in the encoder-decoder model, which allows the model to establish a direct relationship between the input and output sequences. This relationship is captured through the use of attention weights, which assign higher weights to the relevant inputs and lower weights to the non-relevant ones, thereby enabling each output to focus on the appropriate input. By using attention, the decoder can access a weighted combination of all the encoder’s hidden states, rather than relying solely on the last hidden state of the encoder.
In the attention-based encoder-decoder model, at each time step, the output of the decoder is influenced by the previous outputs, $y_{<t}=\left\{y_{1},y_{2},\dots,y_{t-1}\right\}$, the current decoder hidden state, $s_{t}$, and an attention “context” vector, $c_{t}$.
$p(y_{t}|{y_{1},y_{2},\dots,y_{t-1}},x)=f(s_{t},c_{t}),\\ f(s_{t},c_{t})=\text{softmax}(W_{s}\tilde{s}_{t}),\\ \tilde{s}_{t}=\text{tanh}(W_{c}[c_{t},s_{t}]).$
where the $W_{c}$ is a learnable weight matrix. The current hidden state, $s_{t}$, is determined by the decoder GRU unit, which takes the previous hidden state $s_{t-1}$, and the last decoder output value, $y_{t-1}$ as input:
$s_{t}=GRU(S_{t-1},y_{t-1})$
To calculate the context vector, $c_{t}$, the attention model proposed by [1] is used. At time step $t$, the context vector, $c_{t}$, is the weighted average over all the encoder hidden states. The weights, $w_{tt^{\prime}}$, which are the normalized attention scores, $e_{tt^{\prime}}$, determine the amount of attention to each encoder output.
$c_{t}=\sum_{t^{\prime}=1}^{N_{x}}w_{tt^{\prime}}h_{t^{\prime}},\\ w_{tt^{\prime}}=\frac{\exp(e_{tt^{\prime}})}{\sum_{k=1}^{N_{x}}\exp(e_{tk})},\\ e_{tt^{\prime}}=s_{t}^{T}W_{a}h_{t^{\prime}}$
where $e_{tt^{\prime}}$ is called score function in [1], $h_{t^{\prime}}$ is the encoder hidden state at time step $t^{\prime}$, and $W_{a}$ is weight matrix. For a more comprehensive understanding of the attention mechanism, see [56].
V-B Dataset

### v2

# Surrogate Modeling of Trajectory Map-matching in Urban Road Networks using Transformer Sequence-to-Sequence Model

@@ Abstract
Large-scale geolocation telematics data acquired from connected vehicles has the potential to significantly enhance mobility infrastructures and operational systems within smart cities. To effectively utilize this data, it is essential to accurately match the geolocation data to the road segments. However, this matching is often not trivial due to the low sampling rate and errors exacerbated by multipath effects in urban environments. Traditionally, statistical modeling techniques such as Hidden-Markov models incorporating domain knowledge into the matching process have been extensively used for map-matching tasks. However, rule-based map-matching tasks are noise-sensitive and inefficient in processing large-scale trajectory data. Deep learning techniques directly learn the relationship between observed data and road networks from the data, often without the need for hand-crafted rules or domain knowledge. This renders them an efficient approach for map-matching large-scale datasets and more robust to the noise. This paper introduces a deep-learning model, specifically the transformer-based encoder-decoder model, to perform as a surrogate for offline map-matching algorithms. The encoder-decoder architecture initially encodes the series of noisy GPS points into a representation that automatically captures autoregressive behavior and spatial correlations between GPS points. Subsequently, the decoder associates data points with the road network features and thus transforms these representations into a sequence of road segments. The model is trained and evaluated using GPS traces collected in Manhattan, New York. Achieving an accuracy of 75%, transformer-based encoder-decoder models extensively employed in natural language processing presented a promising performance for translating noisy GPS data to the navigated routes in urban road networks.
@@ III-A Map-matching preliminaries
Map-matching is a technique designed to determine the sequence of connected road segments on a digital road network map that best approximates the actual route taken by a vehicle, even in the presence of errors and noise in the observed GPS data. This task is challenging, especially in an urban environment, due to the significant positioning errors and the complex, dense road map. Map-matching task can be categorized into two types: online and offline matching. Online map-matching algorithms handle streaming GPS data in real time, providing immediate feedback on the moving vehicle’s current position relative to the map. In contrast, offline map-matching algorithms process historical trajectory data after it has been fully collected, allowing for more accurate and robust matching decisions by utilizing the entire dataset. This study proposes an offline map-matching algorithm that takes a complete sequence of observed coordinates in raw GPS data as input and produces a segment-based trajectory as the output. However, the low sampling rate in data collection leads to a sparse input representation. Consequently, no direct one-to-one correspondence exists between observed GPS points and inferred road segments.
@@ IV-C Objective Function
Let $g=(g_{1},g_{2},...,g_{n})$ be the transformed version of $T=\left\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{n},y_{n})\right\}$, representing the variable-length input sequence, here in this study, a sequence of grid cells, each of which encompasses each GPS point, and $r=(r_{1},r_{2},...,r_{m})$ be the variable-length output sequence, i.e., sequence of road segments, $m$ is the length of the output sequence. The seq2seq model maps $g$ to $r$ by learning a conditional probability distribution over the set of possible output sequences as:
$p(r|g)=\prod_{t=1}^{m}p(r_{t}|r_{1},\dots,r_{t-1},g)$ (1)
The model learns a probability distribution from input-output pairs in training, adjusting parameters to maximize output sequence likelihood given input; the seq2seq model aims to maximize this log-likelihood, described as:
$\begin{split}arg\max\limits_{\theta}\ P(r|g)=argmax\log P(r|g)\\ =argmax\sum_{t=1}^{m}logP(r_{t}|g_{n},r_{<t}).\end{split}$ (2)
@@ IV-D1 Encoder
$PE_{(pos,i)}=sin(\frac{pos}{10000^{(\frac{2i}{d_{emb}})}}),$ (3)
$PE_{(pos,2i+1)}=cos(\frac{pos}{10000^{(\frac{2i}{d_{emb}})}}),$ (4)
where $pos\in\{1,\dots,n\}$ is the position of the token in the sequence of length $n$, $i\in\{1,\dots,d_{emb}\}$ is the dimension of the positional encoding, $d_{emb}$ is the model’s embedding dimension also referred as model dimension.
@@
$\text{Attention}(Q,K,V)=\text{Softmax}(\frac{QK^{T}}{\sqrt{d_{k}}})V,$ (5)
@@
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V}),$ (7)
where $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$, and $W^{o}$ are learnable weight matrices.
@@ IV-D2 Decoder
V Comparative Evaluation with State-of-the-art Methods
This section presents a comparative evaluation of our proposed model against several state-of-the-art models. The goal of this evaluation is to demonstrate the effectiveness of our model in performing as a surrogate deep learning model for the map matching task. Collecting ground truth data for GPS trajectories is a challenging and resource-intensive task. To the best of authors’ knowledge, the publicly available vehicular trajectory datasets with ground truth labels in our region of interest, New York City, are scarce, making it difficult to perform comprehensive evaluations on real trajectory data. To address this limitation, a data augmentation technique is adopted to generate pairs of noisy trajectories and ground truth trajectories. The following section presents a detailed process of data augmentation that enables easy replication of realistic trajectories.
V-A Data augmentation
In this section, four different combinations of noise level and sampling interval are considered in the data generation. Each combination consists of either low noise (25 meters) or high noise (50 meters) with short intervals (20 seconds) or long intervals (40 seconds). Figure 4 illustrates an example of four different representations of an augmented trajectory with noisy GPS points, each featuring varying noise levels and sampling intervals (SI). For each of the four combinations, 10,000 GPS trajectories are generated for downtown Manhattan road network in New York City shown by Figure 3. Of these, 80% (8,000 trajectories) are randomly selected to train the transformer model, and 20% (2,000 trajectories) are used to test and compare the performances of all three models. Transformer model discards the trajectories with lengths less than four and greater than 200, excluding 16% of the training and testing data in long interval dataset and around 4% in the short interval dataset. The test data is identical across all three models to ensure a fair comparison. The synthesized data for the comparison study component of the paper is available to the public on a permanent data repository at [55]. The data generation process is as follows.
1. A weighted directed graph representing the structure and geospatial features of the road network is created using OpenStreetMap.
2. To generate more realistic and complex paths beyond the shortest route between origin and destination points, since drivers do not always choose the shortest paths, a perturbation technique is applied. The graph’s edge weights are assumed as street lengths except for approximately 30% of the randomly selected edges, the weights are slightly increased to make these streets less preferable.
3. Origin and destination points for each trajectory are randomly selected from the map. The Dijkstra algorithm is then used to find the shortest path between each pair based on the adjusted edge weights on the weighted directed graph. The path is accepted if its length does not exceed a threshold of 6 miles. This constraint is imposed to avoid overly long routes, simplifying model evaluation in this section. In the actual application during the preprocessing phase, a data segmentation technique is employed to handle long trajectories.
4. The generated routes, structured as sequences of road segments, are considered the true paths from which noisy GPS points are collected. These routes serve as the true labels for the map-matching task.
5. To simulate noisy GPS points, first, vehicle speed data and traversed distances are used to generate true GPS points. It is assumed that the vehicle’s speed equals the posted speed limit. Based on the speed and the length of each road segment, true GPS points, representing the vehicle’s actual location, are sampled at specified sampling intervals. Then, Gaussian noise with a specified standard deviation is added to each true GPS point to create the noisy GPS points as the input for the map-matching model.
Figure 3: Street network map of Downtown Manhattan, New York. The base map is provided by OpenStreetMap contributors.
Figure 4: Illustration of an augmented trajectory with different noise level and sampling intervals (SI).
V-B State-of-the-art map-matching methods
Different map-matching models necessitate various input features to execute map-matching. Some models operate solely on noisy GPS points [33], while others require additional heading information [38] or in-vehicle speed information [22]. Due to the absence of speed and heading information in our dataset, and considering that multiple heading changes typically occur within 30-60 second sampling intervals [29]—making heading computation non-trivial without comprehensive contextual data—we have chosen to focus on models that perform map-matching using only sequences of noisy GPS points for comparative analysis.
V-B1 ST-matching [36]
ST-matching is one of the popular map-matching algorithms that has gained extensive attention [56]. This algorithm takes into account the spatial geometric and topological structures of the road network, as well as the temporal constraints of the trajectories. The parameters for this algorithm are set as follows: the maximum number of candidates is $k=10$, the search radius is $r=500$ meters, and the GPS error is $\sigma=25$ meters for low-noise datasets and $\sigma=50$ meters for high-noise datasets.
V-B2 Leuven-matching: HMM with non-emitting states [57]
In conventional Hidden Markov Models (HMMs), each state transition absorbs an observation, leading to issues when measurements are less frequent than segment transitions. A novel map-matching algorithm inspired by bioinformatics is introduced by [57] that employs non-emitting states enabling dynamic interpolation without needing an observation for each transition. The parameters of this algorithm are set as follows: the observation noise is $\sigma=25$ meters for low-noise datasets and $\sigma=50$ meters for high-noise datasets, possible candidate states per observation 10, and minimum normalized probability of observations as 0.0001.
V-B3 Transformer-based surrogate map-matching
The configuration of our deep-learning surrogate model is set as follows: The embedding dimension ($d_{emb}$) is set to 512, while the MLP intermediate dimension ($d_{ff}$) is set to 1024. We use $N=4$ transformer blocks for both the encoder and decoder, and each multi-head attention block employs $h=4$ linear transformations, resulting in $d_{k}=128$. The Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function is utilized. The batch size is set to 32, and the model trained for a total of 200 epochs. The inference time for map-matching using the constructed transformer model varies with trajectory length, with an average inference time of 0.30s and a standard deviation of 0.16s for trajectories ranging from 6 to 204 road segments.
V-C Data structure
The ST-Matching and Leuven-Matching models both are instantiated with a weighted directed graph representing the road network. In this graph, the edges correspond to road segments, and the nodes correspond to the endpoints of these segments. This graph is created using the OSMnx Python package, which reads and represents street features from OpenStreetMap as a graph [58]. These models take as input a sequence of noisy GPS points and output a sequence of road segments, represented by the nodes of endpoints of these segments.
For the transformer model, the same graph generated from OpenStreetMap is utilized. Additionally, a list of geometries is pre-generated to represent the polygons of uniform grid cells, partitioning the region of interest into exclusive equal blocks, each with a unique identifier. More detailed information is provided in Section VI-C. The transformer model is instantiated with these geometries and two vocabulary objects for input and output: the set of all input tokens (in this spatial context, tokens are the grid cells, each assigned a unique index) and the collection of all output tokens (here the road segments from OpenStreetMap, each assigned a unique index), respectively. The input to the transformer model is initially noisy GPS points, which are mapped to the grid cells within which each point lies, and the output is a sequence of the road segments (endpoints of the segments).
V-D Evaluation Metrics
The performance of each model is evaluated using three different metrics, accuracy, Jackkard similarity, and Bilingual Evaluation Understudy (BLEU) [59], mathematically described below:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{T}_{i}\cap T_{i}|}{|T_{i}|}$ (8)
where $T_{i}$ and $\hat{T}_{i}$ are the $i$th ground truth path and the estimated path in the whole trajectory set $T$. |.||.|| . | represents the segment count of the trajectory.
The Jaccard similarity, shown by equation 9 is a measure of similarity between two sets. Here, we use this metric to compare each trajectory with model estimation. It is defined as the size of the intersection of the two routes composed of road segments divided by the size of their union and yields a value between 0 and 1, where 0 indicates no similarity (no common elements) and 1 indicates complete similarity (identical sets).
$\text{Jackkard}=J(T_{i},\hat{T}_{i})=\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{T}_{i}\cap T_{i}|}{|\hat{T}_{i}\cup T_{i}|}$ (9)
BLEU is a metric commonly used to evaluate the quality of machine-generated text, particularly in the context of machine translation. However, in the context of this paper, BLEU is utilized to evaluate the similarity of the predicted segment-based trajectory with the ground truth trajectory. Similar to the text evaluation, here it is assumed to operate by comparing n-grams (contiguous sequences of n road segments) in the candidate and reference trajectories. The intuition behind BLEU is that a good mapping should have similar n-gram statistics to those of the ground truth trajectories. BLEU which is shown in equation 12, has two parts, precision, see equation 10, and brevity penalty, see equation 11. Precision which counts how many contiguous sequences of “n” items (n-grams) in the candidate route match those in the reference route and divides this count by the total number of n-grams in the candidate routes and brevity penalty that penalizes the short mappings [59].
$Precision(\hat{T}_{i},T_{i})=\frac{\text{count of n-gram matches between }\hat{T}_{i},T_{i}}{\text{Count of n-grams in }\hat{T}_{i}},$ (10)
$BP=\begin{cases}1&|\hat{T}_{i}|>|T_{i}|\\ \exp{(1-\frac{|\hat{T}_{i}|}{|T_{i}|})}&|\hat{T}_{i}|\leq|T_{i}|,\\ \end{cases},$ (11)
$BLEU=BP\exp(\frac{1}{n}\sum_{i=1}^{n}\log(Precision_{i}))$ (12)
V-E Comparison results
The map-matching results are presented in table I. Although increasing the noise and the sampling interval degrades the performance of all three methods, the transformer consistently outperforms the other two state-of-the-art methods, ST-Matching and Leuven-Matching, across all four settings. One possible reason for this observation is that the state-of-the-art methods rely on the shortest path as the most probable route between two points. In contrast, the deep learning model not only learns the noise patterns but also captures drivers’ driving patterns and route selection behavior by analyzing large collections of trajectories. An example of a successful capture of the true path using the transformer model is illustrated in Figure 5. While the need for the large training dataset is a limitation of deep learning models, techniques such as transfer learning can address issues related to sparse datasets. Although this is not the focus of this study, training the model with a sparse dataset is an area for future research.
TABLE I: Comparison of Models Across Different Settings
Setting Model Accuracy J BLEU
LN-SI1 Transformer 0.8529 0.9182 0.8451
ST-matching 0.7741 0.8531 0.7991
Leuven-matching 0.7462 0.7978 0.7053
LN-LI2 Transformer 0.8004 0.8728 0.8012
ST-matching 0.6985 0.7931 0.7296
Leuven-matching 0.5972 0.6524 0.5608
HN-SI3 Transformer 0.8386 0.8984 0.8343
ST-matching 0.7170 0.8326 0.7239
Leuven-matching 0.8158 0.9025 0.7862
HN-LI4 Transformer 0.7698 0.8473 0.7725
ST-matching 0.6632 0.7743 0.6915
Leuven-matching 0.5987 0.6734 0.5776
  • 1
    LN-SI: Low Noise - Short Interval
  • 2
    LN-LI: Low Noise - Long Interval
  • 3
    HN-SI: High Noise - Short Interval
  • 4
    HN-LI: High Noise - Long Interval
Figure 5: An example of trajectory map-matching results (right figure) using ST-Matching (green), and Leuven-Matching (orange), Transformer (blue) for an augmented trajectory with the true path represented in black (left figure). Between two initial GPS points, the transformer accurately estimates the true path compared to the other two methods.
VI Experiment with real data
VI-A Baseline
VI-A1 HMM
@@
VI-A2 RNN-based encoder-decoder with attention mechanism
@@
The main components of the RNN-based encoder-decoder architecture are two layers: the embedding layer and the RNN layer. Therefore, the main difference between the RNN-based encoder-decoder with the transformer-based encoder-decoder is its recurrent layer versus the self-attention layer in the transformer. The architecture of the RNN-based encoder-decoder model can be found in [61]. In an RNN-based encoder-decoder, during the encoding, the RNN layer processes all the tokens on the input sequence iteratively and reutilizes the hidden state of an iteration as the input for generating the hidden state for the next element. The last hidden state of the RNN is passed as input to the first iteration of the decoder. Therefore, when processing a sequence of inputs, the RNN layer generates a sequence of hidden states, with each hidden state being a function of the current input and the previous hidden state. In particular, a bidirectional Gated Recurrent Unit (Bi-GRU) in the RNN layer of the encoder, and a regular GRU in the RNN layer of the decoder are utilized. The structure of the Gated Recurrent Unit (GRU) is presented by [62]. GRU captures the sequential information in the data by maintaining an internal hidden state that evolves over time. It has a gating mechanism that controls the flow of information within the network, making it capable of learning and remembering longer sequences compared to traditional RNNs. The gating mechanism consists of an update gate and a reset gate, which control how much of the previous hidden state should be retained and how much new information should be incorporated in each time step. A Bi-GRU is a variant of the GRU architecture that processes input sequences in both forward and backward directions. The input sequence is fed into two separate GRU layers, one processing the input sequence in the forward direction and the other processing the input sequence in the backward direction. The outputs of the two layers at each time step are concatenated and passed to the next layer or used as the final output.
VI-B Dataset